If a + b = 3 and 2b + c = 2, then 2a - c = ?
a) -4
b) -1
c) 1
d) 4
e) 5
Again and again thanks for helping! wonderful website and tutors :-)

Answer by Edwin McCravy(20081)

a + b = 3 and 2b + c = 2, then 2a - c = ?
a + b = 3

Add -a to both sides:

    b = 3 - a

Substitute (3 - a) for b in the second equation: 

       2b + c = 2
 2(3 - a) + c = 2

Use the distributive principle to remove the parentheses:

   6 - 2a + c = 2

Add -6 to both sides:

      -2a + c = -4

Multiply both sides by -1:

  -1(-2a + c) = -1(-4)

Use the distributive principle to remove the parentheses:

       2a - c = 4 

Choice d.

Edwin